Show that x=Asinwt+Acoswt, satisfies the differential equation md^2x/dt^2 + kx=0
1
Expert's answer
2020-02-10T09:25:49-0500

"x=A\sin \omega t+A\cos \omega t"


"\frac{dx}{dt}=A\omega\cos\omega t-A\omega\sin\omega t"

"\frac{d^2x}{dt^2} =\frac{d}{dt}(A\omega\cos\omega t-A\omega\sin\omega t)= -A\omega^2 \sin\omega t-A\omega^2\cos\omega t"


Substitute x into this equation: "m\frac{d^2x}{dt^2}+kx=0"

"m( -A\omega^2 \sin\omega t-A\omega^2\cos\omega t)+k( A\sin\omega t+A\cos\omega t)=0"

"A(k-m\omega ^2)(\sin\omega t+\cos \omega t)=0" for all t


Therefore, x satisfies the differential equation "m\frac{d^2x}{dt^2}+kx=0"

if "A=0" or "k-m\omega ^2=0."